# ClusterRCA: Network Failure Diagnosis in HPC Systems Using Multimodal Data

### v1

@@ Abstract
Index Terms:
Network failure diagnosis, HPC system, Multimodal data, Random walk
@@ I Introduction
High-performance computing (HPC) systems are essential for large-scale applications across various domains like weather forecasting, biological sciences, and fluid dynamics simulation [1]. These systems, typically organized into network clusters, facilitate complex computations and data processing. However, the intricate network structures of HPC systems make them susceptible to issues like network congestion and interconnect failures [2]. Therefore in the domain of HPC, accurate and efficient failure diagnosis is crucial for restoring network performance and ensuring uninterrupted operation. Failure diagnosis involves localization and classification[38], where localization helps narrow down troubleshooting by identifying the primary culprit node, and classification enables prompt determination of mitigation measures based on the specific failure type. Therefore, network failure diagnosis, including localization and classification, is critical for HPC system maintenance.
When a task is observed to be faulty(e.g., running timeout), the operator performs failure diagnosis on the components involved in this task, including root cause localization and type classification. Components in HPC systems typically include: compute nodes, network switches, storage systems, control nodes etc. Among these components, the compute nodes and network switches are mainly responsible for performing the tasks, while network failures mainly occur among them. In contrast, failures on pure interconnected networks (i.e., spine switches, without computing nodes) have been extensively studied and are relatively mature [35] [37]. Therefore in this work, we focus exclusively on network failures localization and classification at computing nodes and connected network switches(i.e., leaf switches). In addition, we use “node” to refer to compute node and network switch later in the paper.
@@
As shown in Fig 1, the HPC system has the following characteristics:
@@
    The compute nodes and network switches are connected via a network interface controller (NIC) in a preconfigured topology. Synchronized communication between computing nodes is extremely dependent on the interconnection network; the presence of network switches as important intermediate network devices connecting them is indispensable to the network composition. This topology makes the occurrence of failures in both correlated.
@@
    In order to monitor the network performance of HPC systems while ensuring that the systems deliver maximum arithmetic performance, the monitoring systems typically collect two types of monitoring data on the compute nodes and network switches during the execution of the application: logs and coarse-grained metrics (minute level). Logs are semi-structured text that describes the executing state of the devices, including state changes, software, and hardware errors, etc. Metrics are multivariate time series, including CPU utilization, network I/O utilization, etc.
@@
    As listed in Table I, different data modalities reflect various types of network failures. Analyzing data from just one modality is insufficient for precise failure diagnosis, as some faults are only reflected in metrics or logs. Besides metrics and logs, operators also document the topology of these devices. For example, due to the NIC sending queue of the computing node being full, network packet transmission timeout occurs, resulting in a large number of PFC messages. Or, the inability of the switching port to adjust the outgoing packet rate of the surrounding compute nodes results in the relevant compute nodes constantly receiving packets that exceed their processing capacity limits, at which point the metric “rx_discards_phy”, which records the number of received packets that are discarded, will rise significantly.
@@
    In our scenario, the monitoring data from compute nodes and network switches is heterogeneous. For example, for the network switch, we only need to record network resource metrics, while for the compute node, we need to record additional CPU resource metrics. Similarly, it can be noted from Table I that network failures will only occur on one of the components, except when the link connecting the two fails (e.g., PFC parameter unaligned).
In recent years, failure localization and classification methods can be categorized into two types: classifier-based [18, 20, 21, 22] and graph-based [23, 24, 19, 25]. Classifier-based methods attempt to directly distinguish the root causes from others by analyzing the features of monitoring data. However, the HPC system oriented approach [18] is only for compute nodes and does not consider the heterogeneous data on the network switches and the correlation between the two. For other scenarios (e.g., cloud and microservices) [20, 21, 22], only a single modality of data is considered, which is limited in terms of the accuracy of the final diagnostic results. Graph-based methods aim to establish causal relationships (i.e., directed edges) between nodes to diagnose the root cause. However, most of these approaches focus on metrics, mining correlations between nodes through network traffic data or known service relationship information. In HPC scenarios, the collection of network metrics is often limited to a coarse-grained level in order to alleviate the computational resources occupied by the monitoring system, which, coupled with the lack of trace data that can directly represent dependencies, greatly increases the difficulty of graphical modeling. Then later, some multimodal methods [3, 39, 40] combine the two. After using the classifier to unify the representation and feature extraction of multimodal data, the graph is then given to perform accurate root cause localization, which solves the problem of accuracy as well as the problem of difficult modeling. Similarly, some methods [39, 40] are applied in microservice scenarios, and the construction of the graph still requires trace or service relationship information. Other method [3] can only output the category of the failure and cannot be localized directly to a compute node or network switch.
In the context of HPC systems, since both classifier-based and graph-based approaches have their own advantages, our idea is to explore the combination of these two methods and their application in HPC systems through multimodal data to achieve more accurate failure diagnosis. However, in this process, we face two major challenges:
@@
    Device and data heterogeneity. Because of the way compute nodes and network switches are connected, it is not possible to separate them independently for failure diagnosis, which would ignore their correlation.
@@
    Combination of classifier and graph. Data from two heterogeneous components make classifier-based methods struggle to extract data features. Moreover, graph construction algorithms perform poorly on coarse-grained data (e.g., PC algorithm[25, 19]) or lack logical dependency data (e.g., trace data[34, 40], network traffic data[23, 24]). Previous methods lack accuracy in single diagnostic targets (i.e., localization and classification) in HPC system scenarios, making it even more difficult to combine the two targets.
To tackle the above challenges, we propose ClusterRCA, which integrates multimodal data for analysis, localizing the culprit node, and determining the failure type of network failures in HPC systems. For the special structure of HPC system, ClusterRCA first translates logs and metrics into events through feature engineering, aggregating devices into NIC pairs based on the topological structure for unified modeling. To accurately localize culprit NIC pairs and determine failure types, ClusterRCA proposes integrating classifier-based and graph-based methods. First, it utilizes the output results of the classifier to establish directed graphs. Then, it adopts graph-based techniques for precise culprit NIC pairs localization and failure type classification by a customized random walk technique. Finally, based on the type of failure, it can determine whether the failure originated from a compute node or a network switch (or the link connecting them), so that the operator can carry out the corresponding remediation measures.
@@
  • We introduce a feature engineering technique based on NIC pairs. Utilizing the NIC pairs in the network topology directly connecting the compute nodes and the network switches as the basic units, we unify the metric and log data features from heterogeneous equipment devices while taking into account the correlation between the two components, to address the challenge of conducting joint analysis posed by the complex structure of HPC systems.
  • To precisely localize the culprit node and determine the failure type, we propose a customized random walk technique with the help of a random forest. We use the classification results from random forest to guide the transitions in random walk, addressing the accuracy limitations inherent in standalone classifier-based and graph-based methods.
  • We conduct extensive experiments using data collected by a top-tier global HPC device vendor, evidencing ClusterRCA’s diagnostic ability.
@@ II-A HPC Systems and Monitor Data
Topology: When an HPC application executes, the computing nodes participating in the application are selected, and the topology information indicates their interconnection edges within the cluster. Typically, a computing node first connects to a switch port and then to other computing nodes. Computing nodes have their own NIC, Switch nodes have multiple NICs for each port, and NICs connect via a physical link. The interconnection between NICs constitutes topology, enabling operators to analyze the potential impact relationships between nodes when network issues arise. Topology is recorded as a dictionary, e.g., (“node”: “server1”, “nic”: “NIC1”, “link”: “switch”, “link_port”: “100GE/18”) indicating that the NIC “NIC1” from the computing node “server1” interconnects with the 18th 100GE port of the switch “switch1”. For each running application, there is a corresponding topology dictionary.
@@
Log: Logs are semi-structured with timestamps and levels [10], recording software and hardware events from devices, such as software and hardware errors from computing nodes, the status of traffic control mechanisms and configuration change events from switches, etc. Operators frequently use the content and quantity information printed in logs to infer device operations and status. To efficiently analyze the vast number of logs generated by various devices, researchers have proposed several methods to automatically parse log information and extract it as templates and parameters[7, 11, 8, 9].
@@ II-B Categories of Network Failures
Network congestion and interrupt significantly impact HPC applications and remain an active area of research[2, 32, 12]. A stable HPC system network depends on the performance of cluster network devices, physical links, and flow control protocols. NICs are virtual network devices in HPC systems that analyze network traffic data. Proper NIC operation is crucial to cluster network performance. NIC failures arise from software and hardware failures, frequently causing incorrect parsing or message sending and inevitably affecting node communication with interconnected nodes. Therefore, NIC failure is one major category of root causes that lead to network failures, characterized by a large volume of error processing packets generated by the node’s NIC, e.g., improper packet errors and packet loss. Secondly, as physical links directly connect to nodes’ NICs, uncontrollable environmental factors can cause link bottlenecks, preventing node traffic from meeting expectations and triggering many flow control mechanisms. Flow control failures refer to unsatisfactory flow control mechanism performance on cluster nodes. Flow control mechanisms are essential to maintain lossless network performance, and failure causes unprocessed flow control packets and packet loss due to excessive traffic. The types of root causes for network issues that we focus on are listed in Table I.
Even when originating at a single node in the cluster, the failures above will propagate expeditiously at a time far shorter than the collection interval of traffic monitoring metrics. Failure symptoms (i.e., packet errors, packet loss, and flow control messages) spread from the culprit node, obscuring the impact of the culprit nodes and victim nodes and generating complex failure patterns.
@@ II-C Problem Statement
When a network failure occurs, operators need to localize the culprit node and determine the failure type to achieve timely failure mitigation. For HPC systems, the first task is a ranking problem: to rank the culprit node (computing node or switch port) higher than other nodes. The second task is a classification problem: to classify a network failure into a well-defined set of failure types. In this paper, we only focus on failures occurring between leaf switches and computing nodes, as achievements have already been made by [35] [37] in the failure localization work between switches.
@@ III-A Design Overview
Figure 2: The overview of ClusterRCA. Solid lines denote processes that contain offline training and online diagnosis, and dash lines denote only online diagnosis processes.
In this paper, we propose a novel framework for localizing culprit nodes and determining failure types upon network failure in HPC systems. As shown in Fig. 2, ClusterRCA comprises two phases: the offline training phase and the online diagnosis phase. The offline training phase will use historical metrics and logs to train feature extraction models and perform feature fusion based on historical topology to obtain the NIC pair feature vectors. In this way, we can represent multimodal data uniformly while ensuring information integrity. Then, the NIC pair feature vectors labeled as culprit, victim, or normal by the operator are input into the state classifier for training with corresponding labels. In this process, the state classifier learns various types of vector patterns in order to gain the ability to discriminate the data. In the online diagnosis phase, real-time metrics, logs, and topology data are used to generate NIC pair feature vectors, which are fed into a trained classifier to obtain the state, and then a state transfer matrix is constructed, which is combined with a state-specific transition strategy and a customized random walk algorithm to accurately localize the culprit node and determine the type of fault.
@@ III-B Feature Engineering
Figure 3: The feature extration and fusion for a NIC pair.
@@ III-B1 Pattern matcher
Metric monitoring of HPC systems often has coarse time granularity to reduce resource consumption, which prevents millisecond-level traffic direction tracking [36]. Therefore, useful information primarily comprises metric change patterns and value levels. PatternMatcher [4] uses a trained neural network model to classify interesting metric change patterns, capturing changes upon congestion or interruption. Fig. 4 (upper) shows the pattern matcher metric pattern extraction process. Inputs to 1D convolution and fully connected networks are normalized time-windowed metric slices. Outputs indicate pattern types. During offline training, operators label historical metric slices with patterns to train the neural networks. We asked an operator to label 1,000 metric slices in 6 hours. To reflect metric mean value levels, as shown in Fig. 4 (lower), we compute the mean of the same metrics across devices at the same time and use box plots [15] to determine levels and mark them with symbols.
@@ III-B2 Log cluster
Some network failures also cause anomaly logs on the culprit and victim nodes. Anomaly log types and counts are critical features to distinguish the root causes. Firstly, we parse semi-structured logs into log templates. Then, we apply hierarchical clustering on the vectors converted from template2vector to obtain different types of log clusters. Finally, we analyze their numbers under normal conditions by computing cluster means and variances and using the 3-sigma principle to represent NIC pair log cluster quantitative features. During offline training, we retain the extracted log template, template2vector model, clustering details, cluster means, and cluster variances as model information for online detection. The log feature extraction process is shown in Fig. 5.

### v2

# ClusterRCA: An End-to-End Approach for Network Fault Localization and Classification for HPC System

@@ I Introduction
High-performance computing (HPC) systems are essential for large-scale applications across various domains, such as weather forecasting, biological sciences, and fluid dynamics simulation [1]. These systems, typically organized into network clusters, facilitate complex computations and data processing. However, the intricate network structures of HPC systems make them susceptible to issues like network congestion and interconnect failures [2]. Therefore, in the domain of HPC, accurate and efficient failure diagnosis is crucial for restoring network performance and ensuring uninterrupted operation. Failure diagnosis involves localization and classification[3], where localization helps narrow down troubleshooting by identifying the primary culprit node, and classification enables prompt determination of mitigation measures based on the specific failure type. Therefore, network failure diagnosis, including localization and classification, is critical for HPC system maintenance.
When a task is observed to be faulty(e.g., running timeout), the operator performs failure diagnosis on the components involved in this task, including root cause localization and type classification. Components in HPC systems typically include: compute nodes, network switches, storage systems, control nodes etc. Among these components, the compute nodes and network switches are mainly responsible for performing the tasks, while network failures mostly occur among them. In contrast, failures on pure interconnected networks (i.e., spine switches, without computing nodes) have been extensively studied and are relatively mature [4] [5]. Therefore, in this work, we focus exclusively on network failures localization and classification at computing nodes and connected network switches(i.e., leaf switches). In addition, we use “node” to refer to a compute node and network switch later in the paper.
@@
As shown in Fig. 1, the HPC system has the following characteristics:
@@
    The compute nodes and network switches are connected via a network interface controller (NIC) in a preconfigured topology. Synchronized communication between computing nodes is highly dependent on the interconnection network; network switches are indispensable to the network composition as necessary intermediate network devices connecting them. This topology correlates with the occurrence of failures in both.
@@
    To monitor the network performance of HPC systems while ensuring that the systems deliver maximum arithmetic performance, the monitoring systems typically collect two types of monitoring data on the compute nodes and network switches during the execution of the application: logs and coarse-grained metrics (minute level). Logs are semi-structured text describing the devices’ executing state, including state changes, software, and hardware errors, etc. Metrics are multivariate time series, including CPU utilization, network I/O utilization, etc. Besides metrics and logs, operators also document the topology of these devices.
@@
    As listed in Table I, different data modalities reflect various types of network failures. Analyzing data from just one modality is insufficient for precise failure diagnosis, as some faults are only reflected in metrics or logs. For example, due to the NIC sending queue of the computing node being full, network packet transmission timeout occurs, resulting in a large number of PFC messages. Or, the inability of the switching port to adjust the outgoing packet rate of the surrounding compute nodes results in the relevant compute nodes constantly receiving packets that exceed their processing capacity limits, at which point the metric “rx_discards_phy”, which records the number of received packets that are discarded, will rise significantly.
@@
    In our scenario, the monitoring data from compute nodes and network switches is heterogeneous. For example, we only need to record network resource metrics for the network switch, while for the compute node, we need to record additional CPU resource metrics. Similarly, it can be noted from Table I that network failures will only occur on one of the components, except when the link connecting the two fails (e.g., PFC parameter unaligned).
In recent years, failure localization and classification methods can be categorized into two types: classifier-based [6, 7, 8, 9] and graph-based [10, 11, 12, 13]. Classifier-based methods attempt to directly distinguish the root causes from others by analyzing the features of monitoring data. However, the HPC system oriented approach [6] is only for compute nodes and does not consider the heterogeneous data on the network switches and the correlation between the two. For other scenarios (e.g., cloud and microservices) [7, 8, 9], only a single modality of data is considered, which is limited in terms of the accuracy of the final diagnostic results. Graph-based methods aim to establish causal relationships (i.e., directed edges) between nodes to diagnose the root cause. However, most of these approaches focus on metrics, mining correlations between nodes through network traffic data or known service relationship information. In HPC scenarios, the collection of network metrics is often limited to a coarse-grained level to alleviate the computational resources occupied by the monitoring system, which, coupled with the lack of trace data that can directly represent dependencies, greatly increases the difficulty of graphical modeling. Then later, some multimodal methods [14, 15, 16] combine the two. After using the classifier to unify the representation and feature extraction of multimodal data, the graph is then given to perform accurate root cause localization, which solves the problem of accuracy and the problem of difficult modeling. Similarly, some methods [15, 16] are applied in microservice scenarios, and the construction of the graph still requires trace or service relationship information. Another method [14] can only output the failure category and cannot be localized directly to a compute node or network switch.
Since both classifier-based and graph-based approaches have advantages in the context of HPC systems, we aim to explore combining these methods and their application in HPC systems through multimodal data to achieve more accurate failure diagnosis. However, in this process, we face two major challenges:
@@
    Device and data heterogeneity. Because of how compute nodes and network switches are connected, it is impossible to separate them independently for failure diagnosis, which would ignore their correlation.
@@
    Combination of classifier and graph. Data from two heterogeneous components makes classifier-based methods struggle to extract data features. Moreover, graph construction algorithms perform poorly on coarse-grained data (e.g., PC algorithm[13, 12]) or lack logical dependency data (e.g., trace data[17, 16], network traffic data[10, 11]). Previous methods lack accuracy in single diagnostic targets (i.e., localization and classification) in HPC system scenarios, making it even more challenging to combine the two targets.
To tackle the above challenges, we propose ClusterRCA, which integrates multimodal data for analysis, localizing the culprit node, and determining the failure type of network failures in HPC systems. For the special structure of the HPC system, ClusterRCA first translates logs and metrics into events through feature engineering, aggregating devices into NIC pairs based on the topological structure for unified modeling. To accurately localize culprit NIC pairs and determine failure types, ClusterRCA proposes integrating classifier-based and graph-based methods. First, it utilizes the output results of the classifier to establish directed graphs. Then, it adopts graph-based techniques for precise culprit NIC pairs localization and failure type classification by a customized random walk technique. Finally, based on the type of failure, it can determine whether the failure originated from a compute node or a network switch (or the link connecting them), so that the operator can carry out the corresponding remediation measures.
@@
  • We introduce a feature engineering technique based on NIC pairs. Utilizing the NIC pairs in the network topology, directly connecting the compute nodes and the network switches as the basic units, we unify the metric and log data features from heterogeneous equipment devices while taking into account the correlation between the two components, to address the challenge of conducting joint analysis posed by the complex structure of HPC systems.
  • To precisely localize the culprit node and determine the failure type, we propose a customized random walk technique with the help of a random forest. We use the classification results from random forest to guide the transitions in random walk, addressing the accuracy limitations inherent in standalone classifier-based and graph-based methods. For greater reproducibility, our sample dataset and code are now publicly accessible. [^1] https://anonymous.4open.science/r/ClusterRCA-F854
  • We conduct extensive experiments using data collected by a top-tier global HPC device vendor, evidencing ClusterRCA’s diagnostic ability. Specifically, results show that ClusterRCA achieves an average top-5 accuracy (Avg@5) of 0.9962, representing an average improvement of approximately 0.378 over the baseline methods, thereby demonstrating its superior network fault diagnosis performance.
@@ II-A HPC Systems and Monitor Data
Topology: In this paper, topology mainly refers to the connection structure between switches and computing nodes in the cluster. When an HPC application executes, the computing nodes participating are selected, and the topology information indicates their interconnection edges within the cluster. Typically, a computing node first connects to a switch port and then to other computing nodes. Computing nodes have their NICs, Switch nodes have multiple NICs for each port, and NICs connect via a physical link. The interconnection between NICs constitutes topology, enabling operators to analyze the potential impact of relationships between nodes when network issues arise. Topology is recorded as a dictionary, e.g., (“node”: “server1”, “nic”: “NIC1”, “link”: “switch1”, “link_port”: “100GE/18”) indicating that the NIC “NIC1” from the computing node “server1” interconnects with the 18th 100GE port of the switch “switch1”. For each running application, there is a corresponding topology dictionary.
@@
Log: Logs are semi-structured with timestamps and levels [20], recording software and hardware events from devices, such as software and hardware errors from computing nodes, the status of traffic control mechanisms, and configuration change events from switches, etc. Operators frequently use the content and quantity information printed in logs to infer device operations and status. To efficiently analyze the vast number of logs generated by various devices, researchers have proposed several methods to automatically parse log information and extract it as templates and parameters[21, 22, 23, 24].
@@ II-B Categories of Network Failures
Network congestion and interruptions significantly impact HPC applications and remain an active area of research[2, 25, 26]. A stable HPC system network depends on the performance of cluster network devices, physical links, and flow control protocols. NICs are virtual network devices in HPC systems that analyze network traffic data. Proper NIC operation is crucial to cluster network performance. NIC failures arise from software and hardware failures, frequently causing incorrect parsing or message sending and inevitably affecting node communication with interconnected nodes. Therefore, NIC failure is one major category of root causes that lead to network failures, characterized by a large volume of error processing packets generated by the node’s NIC, e.g., improper packet errors and packet loss. Secondly, as physical links directly connect to nodes’ NICs, uncontrollable environmental factors can cause link bottlenecks, preventing node traffic from meeting expectations and triggering many flow control mechanisms. Flow control failures refer to unsatisfactory flow control mechanism performance on cluster nodes. Flow control mechanisms are essential to maintain lossless network performance, and failure causes unprocessed flow control packets and packet loss due to excessive traffic. The types of root causes for network issues that we focus on are listed in Table I.
Even when originating at a single node in the cluster, the failures above will propagate expeditiously at a time far shorter than the traffic monitoring metrics collection interval. Failure symptoms (i.e., packet errors, packet loss, and flow control messages) spread from the culprit node, obscuring the impact of the culprit nodes and victim nodes and generating complex failure patterns.
@@ II-C Problem Statement
When a network failure occurs, operators need to localize the culprit node and determine the failure type to achieve timely failure mitigation. For HPC systems, the first task is a ranking problem: to rank the culprit node (computing node or switch port) higher than other nodes. The second task is a classification problem: classifying a network failure into a well-defined set of failure types.
@@ III-A Design Overview
Figure 2: The overview of ClusterRCA. Solid lines denote processes that contain offline training and online diagnosis, and dashed lines denote only online diagnosis processes.
This paper proposes a novel framework for localizing culprit nodes and determining failure types upon network failure in HPC systems. As shown in Fig. 2, ClusterRCA comprises two phases: offline training and online diagnosis. The offline training phase will use historical metrics and logs to train feature extraction models and perform feature fusion based on historical topology to obtain the NIC pair feature vectors. This way, we can uniformly represent multimodal data while ensuring information integrity. Then, the NIC pair feature vectors labeled as culprit, victim, or normal by the operator are input into the state classifier for training with corresponding labels. In this process, the state classifier learns various types of vector patterns to gain the ability to discriminate the data. In the online diagnosis phase, real-time metrics, logs, and topology data are used to generate NIC pair feature vectors, which are fed into a trained classifier to obtain the state. Then, a state transfer matrix is constructed, which is combined with a state-specific transition strategy and a customized random walk algorithm to accurately localize the culprit node and determine the type of fault.
@@ III-B Feature Engineering
Figure 3: The feature extraction and fusion for a NIC pair.
@@ III-B1 Pattern matcher
Metric monitoring of HPC systems often has coarse time granularity to reduce resource consumption, which prevents millisecond-level traffic direction tracking [27]. Therefore, valuable information primarily comprises metric change patterns and value levels. PatternMatcher [28] mimics the visual perception of operators observing indicator curves and trains a one-dimensional neural network model to classify patterns of indicator changes that deserve attention (e.g., steady rise, multiple peaks, sharp decline, etc). Ultimately, each category will be mapped to a specific number. Fig. 4 (upper) shows the pattern matcher metric pattern extraction process. Inputs to 1D convolution and fully connected networks are normalized time-windowed metric slices. Outputs indicate pattern types. During offline training, operators label historical metric slices with patterns to train the neural networks. We asked an operator to label 1,000 metric slices in 6 hours. To reflect metric mean value levels, as shown in Fig. 4 (lower), we compute the mean of the same metrics across devices at the same time and use box plots [29] to determine levels and mark them with symbols.
@@ III-B2 Log cluster
Some network failures also cause anomaly logs on the culprit and victim nodes. Anomaly log types and counts are critical features for distinguishing the root causes. The log feature extraction process is shown in Fig. 5. Firstly, we parse semi-structured logs into log templates using the Drain algorithm [21]. Next, these templates are converted into numerical vectors using Template2Vec [30] to effectively represent the template’s semantic information numerically. To reduce dimensionality and group semantically similar events, we then apply the well-established AGNES (Agglomerative Nesting) algorithm [31] to these template vectors. Finally, we analyze their numbers under normal conditions by computing cluster means and variances and using the 3-sigma principle to represent NIC pair log cluster quantitative features. In the offline phase, we retain the extracted log template, Template2Vec model, clustering details, cluster means, and cluster variances as model information for the online phase.
Finally, we will concatenate the metric features and log features extracted from the same NIC pair. The feature data input into the subsequent process for fault diagnosis will include metric pattern features, numerical features, and log clustering quantitative features.